OpenStudy (anonymous):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle?

OpenStudy (anonymous):

First you times 8 by 2 since there are two lengths or two sides equal to each other in a rectangle. then you would subtract 16 from 52 and get 36, now you would divide this by four because you already subtracted the extra length that makes it a triangle. You would get 9. 9 + 8 = 17 and that's your answer :)

jimthompson5910 (jim_thompson5910):

let w be the width

jimthompson5910 (jim_thompson5910):

since The length of a rectangle is 8 in. more than its width, we know 

jimthompson5910 (jim_thompson5910):

All 4 sides have to add to the perimeter 52, so s1 + s2 + s3 + s4 = (w)+(w+8)+(w)+(w+8) = 4w+16 = 52 So you need to solve 4w + 16 = 52